### v1

@@ Abstract
Liquid electrolytes are critical components of next-generation energy storage systems, enabling fast ion transport, minimizing interfacial resistance, and ensuring electrochemical stability for long-term battery performance. However, measuring electrolyte properties and designing formulations remain experimentally and computationally expensive. In this work, we present a unified framework for designing liquid electrolyte formulation, integrating a forward predictive model with an inverse generative approach. Leveraging both computational and experimental data collected from literature and extensive molecular simulations, we train a predictive model capable of accurately estimating electrolyte properties from ionic conductivity to solvation structure. Our physics-informed architecture preserves permutation invariance and incorporates empirical dependencies on temperature and salt concentration, making it broadly applicable to property prediction tasks across molecular mixtures. Furthermore, we introduce—to the best of our knowledge—the first generative machine learning framework for molecular mixture design, demonstrated on electrolyte systems. This framework supports multi-condition–constrained generation, addressing the inherently multi-objective nature of materials design. This unified framework advances data-driven electrolyte design and can be readily extended to other complex chemical systems beyond electrolytes.
@@ Results
[…]
@@
We tested our CGD approach on two different base formulation constraints: (1) EC $>$ 20%; (2) EC $>$ 20% & DMC (dimethyl carbonate) $>$ 20% & EMC (ethyl methyl carbonate) $>$ 20%. For conditional generation, the target conductivity was set to 10 mS/cm and the anion ratio to 0.25, and more than 10,000 samples are generated for evaluation. As shown in Fig. 3f, our CGD method significantly improves the success rate of electrolyte formulation generation compared to the case without classifier guidance. For case (1), the success rate is improved by a factor of 30, while for case (2), it is enhanced by at least three orders of magnitude (without CGD method, there is no single formulation meets the condition among more than 10,000 generated samples). The success rate accounts not only for satisfying the base formulation constraints, but also for generating electrolytes with conductivity and anion ratio values close to the specified targets. These results demonstrate that CGD can effectively guide the generative process toward realistic, multi-constraint electrolyte formulations with high fidelity.
@@ Methods
To train a conditional diffusion model, we first used our predictive model to synthetically create 30,000 different electrolyte formulations with conductivity and anion ratio labels. The synthetic dataset offers a more balanced coverage of the electrolyte design space compared to the experimental dataset and provides improved accuracy in conductivity evaluation compared to the computational dataset. All 62 solvents and 2 Li salts ($\text{LiPF}_{6}$ and LiTFSI) from experiments are randomly sampled to compose different formulations. The molar ratios of solvents and salts sum up to 1 with fixed salt concentration (10 mol %) and temperature (25°C). We select 2 Li salts as they are most commonly used in today’s commercialized lithium-ion battery. As a result, our generative modeling primarily focuses on the solvent design space, which is typically the main target in electrolyte design. Details about the synthetic dataset are included in the Supporting Information.
@@
Supplementary information Additional explanations and details regarding the datasets, model architectures, training and evalution approaches, and results can be found in the supplementary information.
@@
Data and code availability
The computational, experimental and literature datasets, both predictive and generative models to reproduce the results in the paper, as well as the source codes, are provided in the referenced link (https://github.com/bytedance/bamboo).
Figure 1: A predictive and generative electrolyte design workflow reported in this work. a Forward (prediction) and inverse (generation) processes of electrolyte formulation are designed as three-stage workflows, using molecular embeddings (representations of individual component molecules in an electrolyte formulation) and electrolyte embeddings (permutation-invariant representations of entire electrolyte formulation). b Three stages of predictive model for conductivity and anion ratio predictions: (1) A GNN model is trained on single-molecule dataset on multi-property prediction, generating universal molecular embeddings. (2) MD data of around 100,000 different electrolyte formulations are utilized to further construct an informative electrolyte embedding from molecular embeddings. (3) An empirical relation is integrated into the model architecture and fine-tuned with 10,000+ experimental literature conductivity data points. c Three stages of generative model given property conditions: (1) A conditional diffusion model generates electrolyte embeddings based on specified properties. (2) The generated electrolyte embeddings are converted back to molecular embeddings with a decoder. (3) Finally, molecular embeddings are matched with our chemical database to obtain the electrolyte formulation.
Figure 2: Prediction performance a Comparison of model performance across various molecular properties during molecular pretraining. Our GNN model utilizes atom and bond features as input and is based on an EGT model (red bar, “Atom/bond feature + EGT”). Other two models are considered as baseline results (blue bars, “Morgan fingerprint + NN” and “Atom/bond feature + GAT”). b Anion ratio and ionic conductivity obtained from MD simulations using an OPLS force field. The density of data below 0.005 is omitted in the figure. Here, conductivity calculated using Mistry’s method is used given that it aligns generally better with experiments (Fig. S5). c Predictions of our model versus ground truth after experimental fine-tuning for both anion ratio and conductivity. d Example predictions of temperature and concentration dependence of conductivity across various electrolyte systems with empirical equation. $T_{0}$ is a learnable temperature parameters in the empirical equation, which is generally related to glass transition temperature of electrolytes.
Figure 3: Generation performances a Extrapolation of target properties using generative model. To extrapolate conductivity, the anion ratio is fixed at the dataset’s mean value. Similarly, to extrapolate the anion ratio, conductivity is set to its mean. b Example distribution of generated electrolyte formulations given target conductivity (10.0 mS/cm) and anion ratio (0.3). More results of different conditions can be found in Fig. S8. c One example of generated electrolyte formulation from panel Fig. 3b. More examples are listed in Fig. S9 - Fig. S13. d Evaluation of generation using three metrics including MAPE and two diversity scores. e Schematic of classifier-guided conditional diffusion to ensure that the generated formulation satisfies the base formulation constraints. f Performance comparison between conditional generation and classifier-guided generation in satisfying the base formulation constraints.
Supporting Information
@@ B.1 Single-molecule dataset
After data preprocessing and screening, the final dataset consists of 241,414 entries with 11 molecular properties: $\bm{T_{m}}$: melting point, $\bm{T_{b}}$: boiling point, $\bm{n_{D}}$/$\bm{n_{D}^{\textbf{liquid}}}$: (liquid) refractive index, $\bm{pK_{a}}$: acid dissociation constant, $\bm{pK_{b}}$: base dissociation constant, $\bm{\varepsilon}$: dielectric constant, $\bm{\gamma_{s}}$: surface tension, $\bm{\rho}$: mass density, $\bm{\eta}$: viscosity and $\bm{P_{\text{vap}}}$: vapor pressure. Details regarding the final dataset, along with the corresponding data sources and the number of entries for each property label, are provided in Table S3.
@@ B.2 Multi-task learning
[…]
@@
[…]
@@
[…]
@@ C.2 Experimental dataset
The experimental dataset of electrolyte conductivity is obtained by gathering data from public datasets and academic publications. There are two public datasets we collect data from: (1) Polymer Electrolyte Dataset [9]; (2) CALiSol-23 Dataset [8]. For the Polymer Electrolyte Dataset, we first filtered out all polymer electrolytes based on their SMILES representation, as polymer SMILES contain the symbol $\ast$. For the remaining data, we manually reviewed the source publications, examining each paper individually to remove inconsistent data (e.g. solid electrolyte, duplicate but inconsistent entries, …) and correct any errors (e.g. unnormalized molar/weight ratio, incorrect SMILES, …). We also removed the conductivity at extremely low temperature (below phase transition temperature) and concentration (salt molar ratio $<$ 0.0001). For CALiSol-23 Dataset, a similar data processing approach was applied. In addition to these two public datasets, we incorporate conductivity data from 21 more recent studies, with a focus on interfacial stability, particularly in fluorinated electrolyte systems [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 49, 97, 98, 99, 100, 101].
@@ D.2 Empirical relations
[…]
@@ E.3 Electrolyte decoder
To generate high-dimensional molecular embeddings for solvents and salts from the low-dimensional electrolyte embedding, we utilize a learnable seed vectors $\bm{S}\in\mathbb{R}^{N\times d}$ ($N$ is the number of solvents and salts, $d$ is the dimension of molecular embedding) [72] to extract molecular embedding at each position given the electrolyte embedding as the context. The steps are as follows:
@@ E.4 Evaluation metrics
formulation $\text{formulation diversity}=\frac{1}{N}\sum_{i=1}^{N}\min_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\sum_{k=1}^{d}|\bm{v}_{BoM,i}^{(k)}-\bm{v}_{BoM,j}^{(k)}|$ (53)
@@
molecular $\text{molecular diversity}=-\sum_{i=1}^{N}p_{i}\log(p_{i})$ (54)
@@ E.5 Classifier-guided diffusion
[…]
@@
Finally, we utilize a negative ReLU function to ensure proper gradient behavior during sampling. Above all, the gradient should be zero when the condition is satisfied. In other words, one side of the activation function should be constant as in ReLU function. Second, when the condition is not met, the gradient should be positive; thus, a negative version of ReLU is used. Third, since the ReLU function maintains a constant gradient when the condition is unsatisfied, we do not need to adjust the scale of molar ratios to avoid gradient vanishing or exploding issues. The gradient strength can be easily adjusted via the gradient scale, offering flexibility in guidance control. The gradient scales $\bm{s}$ are chosen between 0, 1, 10, 100, 1000 and 10000. When gradient scale is 0, the classifier does not impose any constraint to the generation which falls back to conditional generation with only conductivity and anion ratio targets. We found that for both base formulation constraints we tested in this study — “EC $>$ 20%” and “EC $>$ 20%, DMC $>$ 20% and EMC $>$ 20%”, the success rates are highest when $\bm{s}=1000$.
@@
Data split no F/with F no S/with S no P/with P high viscosity/low viscosity
@@
Table S3: Single-molecule dataset. The table lists the data resources of different molecular properties.
00footnotetext: $\bm{T_{m}}$: melting point; $\bm{T_{b}}$: boiling point; $\bm{n_{D}}$/$\bm{n_{D}^{\text{liquid}}}$: (liquid) refractive index; $\bm{\varepsilon}$: dielectric constant; $\bm{\gamma_{s}}$: surface tension, $\bm{\eta}$: viscosity; $\bm{P_{\text{vap}}}$: vapor pressure.00footnotetext: Total number of data: $\bm{T_{m}}$: 184921; $\bm{T_{b}}$: 7464; $\bm{n_{D}}$: 4252; $\bm{n_{D}^{\text{liquid}}}$: 2592; $\bm{pK_{a}}$: 2568; $\bm{pK_{b}}$: 3283, $\bm{\varepsilon}$: 1286; $\bm{\gamma_{s}}$: 14317; $\bm{\rho}$: 5181; $\bm{\eta}$: 3405; $\bm{P_{\text{vap}}}$: 32184.
@@
Figure S1: Detailed model architectures within the workflow reported in this work. a Overall workflow of both predictive and generative process. Solvents and Li salts are considered separately in the workflow. b Module architecture of GNN model which takes SMILES of a single molecule as input and generates a universal molecular embedding by multi-task learning on 11 molecular properties. c. A self-attention-based aggregation block to merge multiple molecular embeddings into a mixture embedding and ensure permutation invariance. $\times$ is row-wise multiplication, $\odot$ stands for matrix multiplication and $\langle$,$\rangle$ , ⟩ represents inner product. d The empirical relation block for conductivity and anion ratio prediction. For conductivity, there are six empirical learnable parameters (except viscosity) based on the electrolyte embedding, while anion ratio is predicted directly from the electrolyte embedding concatenated with temperature and concentration using a readout MLP layer. Activation layers are omitted from the schematic. e The decoder module which recovers molecular embeddings from mixture embeddings for both solvents and salts, which is further matched to molecules in our electrolyte molecule database.

### v4

@@ Abstract
Liquid electrolytes are critical components of next-generation energy storage systems, enabling fast ion transport, minimizing interfacial resistance, and ensuring electrochemical stability for long-term battery performance. However, measuring electrolyte properties and designing formulations remain experimentally and computationally expensive. In this work, we present a unified framework for designing liquid electrolyte formulation, integrating a forward predictive model with an inverse generative approach. Leveraging both computational and experimental data collected from literature and extensive molecular simulations, we train a predictive model capable of accurately estimating electrolyte properties from ionic conductivity to solvation structure. Our physics-informed architecture preserves permutation invariance and incorporates empirical dependencies on temperature and salt concentration, making it broadly applicable to property prediction tasks across molecular mixtures. Furthermore, we introduce—to the best of our knowledge—the first generative machine learning framework for molecular mixture design, demonstrated on electrolyte systems. This framework supports multi-condition–constrained generation, addressing the inherently multi-objective nature of materials design. As a proof of concept, we experimentally identified three liquid electrolytes with both high ionic conductivity and anion-concentrated solvation structure. This unified framework advances data-driven electrolyte design and can be readily extended to other complex chemical systems beyond electrolytes.
@@ Introduction
  • We further tested generated electrolyte formulations with experimental conductivity measurement and Raman spectroscopy. Among these promising candidates, we identified three formulations that exhibit both high ionic conductivity and anion-concentrated solvation structure.
@@ Results
To obtain a comprehensive and efficient representation of diverse molecules within the electrolyte formulations, we pretrained a GNN model to learn a universal “fingerprint” by predicting various molecular properties with more than 200,000 molecular data. These properties include 11 distinct entries: melting point ($T_{m}$), boiling point ($T_{b}$), (liquid) refractive index ($n_{D}$$n_{D}^{\text{liquid}}$), $K_{a}, p$$K_{b}$, dielectric constant ($\varepsilon$), surface tension ($\gamma_{s}$), density ($\rho$), viscosity ($\eta$) and vapor pressure ($P_{vap}$). The GNN model takes atomic and bond features derived from the chemical SMILES representation [34] as input and outputs a molecular embedding using a modified Edge-augmented Graph Transformer (EGT) model [35, 36]. This embedding is then processed through separate readout blocks with multiple multilayer perceptrons (MLPs), each dedicated to one respective molecular properties.
@@
We tested our CGD approach on two different base formulation constraints: (1) EC >> 20%; (2) EC >> 20% & DMC (dimethyl carbonate) >> 20% & EMC (ethyl methyl carbonate) >> 20%. For conditional generation, the target conductivity was set to 10 mS/cm and the anion ratio to 0.25, and more than 10,000 samples are generated for evaluation. As shown in Fig. 3f, our CGD method significantly improves the success rate of electrolyte formulation generation compared to the case without classifier guidance. For case (1), the success rate is improved by a factor of 30, while for case (2), it is enhanced by at least three orders of magnitude (without CGD method, there is no single formulation meets the condition among more than 10,000 generated samples). The success rate accounts not only for satisfying the base formulation constraints, but also for generating electrolytes with conductivity and anion ratio values close to the specified targets. These results demonstrate that CGD can effectively guide the generative process toward realistic, multi-constraint electrolyte formulations with high fidelity.
Experimental validation
As a proof of concept, we further conducted experiments to validate the properties of the generated formulations. Specifically, we measured ionic conductivity and used Raman spectroscopy to probe the solvation structure of electrolytes. Although experimentally quantifying ratio of anion coordination is not straightforward, Raman spectroscopy can provide insights into the solvation structure by probing the degree of coordination between anion and cation in the solution [52].
To select formulations for experimental validation, we first applied our conditional generative workflow to propose candidates targeting conductivities of 15 or 20 mS/cm together with anion ratios of 0.3 or 0.5. These generated formulations were then screened using our predictive models to refine the selection. Those that have minimal deviation from the target properties were selected. Given the availability of molecules from commercial chemical vendors, expert inspection for stability and solubility, experimental costs and data validity, we selected 18 formulations for experimental validation including 7 reference formulations with commonly seen single solvent (PC, EC, THF, EA, MA, DEC, DMC) and 11 generated formulations with multi-solvents.
Among the 11 generated formulations, 10 mixture formulations were experimentally validated with ionic conductivity within the target range, significantly outperforming the average conductivity of randomly selected experimental formulations (Fig. 4a). More importantly, Raman spectroscopy analysis reveals a clear rightward shift of the characteristic FSI- peak (typically observed between 710 and 730 cm-1) in three high-conductivity generated formulations (around 15 mS/cm) compared to typical single-solvent electrolytes (Fig. 4b). Among these single-solvent electrolytes, DEC and DMC are well-studied for promoting anion-concentrated solvation structures [53, 52]. These spectral shifts suggest a high degree of anion-cation coordination, indicating that in these generated high-conductivity electrolytes FSI- anions are more likely to enter the first solvation shell of Li+. This behavior corresponds to a higher anion ratio and might lead to enhanced interfacial stability. Notably, the formulation “MA/DEC/EA/EMC/LiFSI” exhibits a right shift of the FSI- Raman peak compared to each individual solvent component (EMC and DMC have similar FSI- peak positions [54]), suggesting a solvation structure beyond interpolation of each individual component within the electrolyte. More details about the compositions of solvents and salts within each formulation can be found in Table S6 and Table S7. These experimental results demonstrate that our generative workflow can propose novel and intriguing electrolyte formulations that exhibit both high conductivity and potentially high interfacial stability.
@@ Methods
To train a conditional diffusion model, we first used our predictive model to synthetically create 30,000 different electrolyte formulations with conductivity and anion ratio labels. The synthetic dataset offers a more balanced coverage of the electrolyte design space compared to the experimental dataset and provides improved accuracy in conductivity evaluation compared to the computational dataset. All 62 solvents and 2 Li salts ($\text{LiPF}_{6}$ and LiFSI) from experiments are randomly sampled to compose different formulations. The molar ratios of solvents and salts sum up to 1 with fixed salt concentration (10 mol %) and temperature (25°C). We select 2 Li salts as they are most commonly used in today’s commercialized lithium-ion battery. As a result, our generative modeling primarily focuses on the solvent design space, which is typically the main target in electrolyte design. Details about the synthetic dataset are included in the Supporting Information.
@@
Conductivity measurement
All the experiments in this work were conducted by HEFEI FANGTIYA LLC. (https://m.zhipin.com/companys/0507e6317c04b3f803x73tq8ElM~.html). The solvents used to prepare liquid electrolytes were purchased from various vendors including Shinghwa Advanced Material Group Co., Ltd. (EC, PC, MA, DEC, EA, EP, DME), Aladdin Scientific (THF, DEM, METHF, EMP, DMM, EFA, MTFA), Xiamen Shouneng Technology Co., Ltd. (DOL, EGDEE), Guangzhou Tinci Materials Technology Co., Ltd. (FEC), Hi-Tech Spring (DMC), Hunan Zhonglan Network Technology Co., Ltd. (LiPF6), and Do-Fluoride New Materials Co., Ltd. (LiFSI). All materials were mixed at room temperature (25 °C), and the conductivities were measured using a conductivity meter (FE32-Standard, Mettler-Toledo Group, Switzerland). Each formulation was prepared and measured three times to ensure consistency and reliability.
Raman spectroscopy
Raman measurements were conducted using a DXR2 Raman microscope. A 532 nm laser was used for most samples, while a 785 nm laser was applied for electrolytes that potentially exhibited fluorescence background. The spectral range covered wavenumbers from 100 to 4250 cm-1. All the experiments were performed by Hefei Fangtiya New Material Technology Co., Ltd.. From Figure 4b, we can see that the peak positions of the single solvent-LiFSI electrolyte have the trend of DMC >> DEC >> MA >> EA >> THF >> EC >> PC, which aligns well with Ref. [54] and demonstrates the validity of our Raman spectroscopy measurement.
@@
Supplementary information Additional explanations and details regarding the datasets, model architectures, training and evaluation approaches, and results can be found in the supplementary information.
@@
Data availability
The computational and experimental literature datasets, together with checkpoints for both predictive and generative models are available on HuggingFace (https://huggingface.co/ByteDance-Seed/bamboo_mixer).
Code availability
The source codes of both predictive and generative workflow to reproduce the results from this work are public on Github (https://github.com/ByteDance-Seed/bamboo_mixer).
[…]
@@ B.1 Single-molecule dataset
After data preprocessing and screening, the final dataset consists of 241,414 entries with 11 molecular properties: $\bm{T_{m}}$: melting point, $\bm{T_{b}}$: boiling point, $\bm{n_{D}}$$\bm{n_{D}^{\textbf{liquid}}}$: (liquid) refractive index, $\bm{pK_{a}}$: acid dissociation constant, $\bm{pK_{b}}$: base dissociation constant, $\bm{\varepsilon}$: dielectric constant, $\bm{\gamma_{s}}$: surface tension, $\bm{\rho}$: mass density, $\bm{\eta}$: viscosity and $\bm{P_{\text{vap}}}$: vapor pressure. Details regarding the final dataset, along with the corresponding data sources and the number of entries for each property label, are provided in Table S3.
@@ B.2 Multi-task learning
$A^{\gamma_{s}}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (1)
$B^{\gamma_{s}}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (2)
@@
$\text{ln}(\eta_{0})=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (4)
$A^{\eta}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (5)
$T_{0}^{\eta}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (6)
@@
$A^{P_{\text{vap}}}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (8)
$B^{P_{\text{vap}}}=\text{MLP}(\bm{h}_{m},\text{act=softplus})$ (9)
@@ C.2 Experimental dataset
The experimental dataset of electrolyte conductivity is obtained by gathering data from public datasets and academic publications. There are two public datasets we collect data from: (1) Polymer Electrolyte Dataset [9]; (2) CALiSol-23 Dataset [8]. For the Polymer Electrolyte Dataset, we first filtered out all polymer electrolytes based on their SMILES representation, as polymer SMILES contain the symbol $\ast$. For the remaining data, we manually reviewed the source publications, examining each paper individually to remove inconsistent data (e.g. solid electrolyte, duplicate but inconsistent entries, …) and correct any errors (e.g. unnormalized molar/weight ratio, incorrect SMILES, …). We also removed the conductivity at extremely low temperature (below phase transition temperature) and concentration (salt molar ratio << 0.0001). For CALiSol-23 Dataset, a similar data processing approach was applied. In addition to these two public datasets, we incorporate conductivity data from 21 more recent studies, with a focus on interfacial stability, particularly in fluorinated electrolyte systems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 53, 21, 22, 23, 49, 24, 25, 26, 27, 28].
@@ D.2 Empirical relations
$\text{ln}(A)=\text{MLP}(\bm{h}_{e},\text{act=softplus})$ (28)
$B=\text{MLP}(\bm{h}_{e},\text{act=softplus})$ (29)
$D=\text{MLP}(\bm{h}_{e},\text{act=softplus})$ (30)
$n_{1}=\text{MLP}(\bm{h}_{e},\text{act=softplus})$ (31)
$n_{2}=\text{MLP}(\bm{h}_{e},\text{act=softplus})$ (32)
$T_{0}=\text{MLP}(\bm{h}_{e},\text{act=sigmoid})$ (33)
@@ E.3 Electrolyte decoder
To generate high-dimensional molecular embeddings for solvents and salts from the low-dimensional electrolyte embedding, we utilize a learnable seed vectors $\bm{S}\in\mathbb{R}^{N\times d}$ (NN is the number of solvents and salts, $d$ is the dimension of molecular embedding) [75] to extract molecular embedding at each position given the electrolyte embedding as the context. The steps are as follows:
@@ E.4 Evaluation metrics
$\text{formulation diversity}=\frac{1}{N}\sum_{i=1}^{N}\min_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\sum_{k=1}^{d}|\bm{v}_{BoM,i}^{(k)}-\bm{v}_{BoM,j}^{(k)}|$ (53)
@@
$\text{molecular diversity}=-\sum_{i=1}^{N}p_{i}\log(p_{i})$ (54)
@@ E.5 Classifier-guided diffusion
$\bm{m}=[0,2],\bm{b}=[0.2,0.4],\bm{\tau}=[1,-1]\Leftrightarrow\text{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace Mol 0\textquotedblright\ltxml@oqmark@close{}}>20\%\text{ and \ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace Mol 2\textquotedblright\ltxml@oqmark@close{}}<40\%$ (58)
@@
Finally, we utilize a negative ReLU function to ensure proper gradient behavior during sampling. Above all, the gradient should be zero when the condition is satisfied. In other words, one side of the activation function should be constant as in ReLU function. Second, when the condition is not met, the gradient should be positive; thus, a negative version of ReLU is used. Third, since the ReLU function maintains a constant gradient when the condition is unsatisfied, we do not need to adjust the scale of molar ratios to avoid gradient vanishing or exploding issues. The gradient strength can be easily adjusted via the gradient scale, offering flexibility in guidance control. The gradient scales $\bm{s}$ are chosen between 0, 1, 10, 100, 1000 and 10000. When gradient scale is 0, the classifier does not impose any constraint to the generation which falls back to conditional generation with only conductivity and anion ratio targets. We found that for both base formulation constraints we tested in this study — “EC >> 20%” and “EC >> 20%, DMC >> 20% and EMC >> 20%”, the success rates are highest when $\bm{s}=1000$.
@@
Appendix F Experimental validation
The 18 formulations we used for evaluation are selected out of 29 formulations measured by experiments, including 8 reference single-solvent formulation and 21 generated multi-solvent formulations. During the experiments, we observed that the reference DOL/LiFSI binary electrolyte and 10 generated multi-component electrolytes either did not exhibit distinct peaks in their Raman spectra or displayed unusually high viscosity. These data were considered unreliable and excluded from further evaluation. As for the potential causes of these unreliable data, the DOL/LiFSI binary solution is known to undergo spontaneous polymerization reactions [39]. The absence of distinct Raman features or ultrahigh viscosity in the 10 generated electrolytes may similarly result from polymerization, side reactions, or salt crystallization.
Chemical stability, reactivity, and solubility of electrolytes are not currently accounted for in the model’s predictions and generative processes, and they are inherently complicated. For example, the polymerization of DOL can be suppressed in the presence of certain chemical species, allowing the solution to remain stable [39]. Furthermore, relevant studies and experimental data remain limited. As a potential path forward, future work could incorporate base formulation constraints into our generative workflow, like we did in this work, to embed domain knowledge and better address issues related.
@@ Appendix F Experimental validation
Data split no F/with F no P/with P no S/with S high viscosity/low viscosity
@@
Table S3: Single-molecule dataset. The table lists the data resources of different molecular properties.00footnotetext: $\bm{T_{m}}$: melting point; $\bm{T_{b}}$: boiling point; $\bm{n_{D}}$$\bm{n_{D}^{\text{liquid}}}$: (liquid) refractive index; $\bm{\varepsilon}$: dielectric constant; $\bm{\gamma_{s}}$: surface tension, $\bm{\eta}$: viscosity; $\bm{P_{\text{vap}}}$: vapor pressure.00footnotetext: Total number of data: $\bm{T_{m}}$: 184921; $\bm{T_{b}}$: 7464; $\bm{n_{D}}$: 4252; $\bm{n_{D}^{\text{liquid}}}$: 2592; $\bm{pK_{a}}$: 2568; $\bm{pK_{b}}$: 3283, $\bm{\varepsilon}$: 1286; $\bm{\gamma_{s}}$: 14317; $\bm{\rho}$: 5181; $\bm{\eta}$: 3405; $\bm{P_{\text{vap}}}$: 32184.
@@
Molecule index Molecule abbreviation Molecule full name Molecule SMILES
1 EC ethylene carbonate C1COC(=O)O1
2 PC propylene carbonate CC1COC(=O)O1
3 MA methyl acetate CC(=O)OC
4 EA ethyl acetate CCOC(=O)C
5 DEC diethyl carbonate CCOC(=O)OCC
6 DOL 1,3-dioxolane C1COCO1
7 THF tetrahydrofuran C1CCOC1
8 METHF 2-Methyltetrahydrofuran CC1CCCO1
9 DMM dimethoxymethane COCOC
10 DEM diethoxymethane CCOCOCC
11 FEC fluoroethylene carbonate C1C(OC(=O)O1)F
12 MTFA methyl trifluoroacetate COC(=O)C(F)(F)F
13 EMC ethyl methyl carbonate CCOC(=O)OC
14 EFA ethyl fluoroacetate CCOC(=O)CF
15 DME 1,2-dimethoxyethane COCCOC
16 EGDEE 1,2-diethoxyethane CCOCCOCC
17 EP ethyl propionate CCC(=O)OCC
Table S6: Solvent molecule list used in this work. The list provides the molecular information of solvents used in formulations of this work.
Formulation
index
Components Molar ratios
Conductivity
(mS/cm)
Peak wavenumber
(cm-1)
1 EC/MA/EA/DOL/DEM/FEC/LiPF6/LiFSI 9%/4%/18%/13%/41%/6%/2%/7% 18.77 713.1
2 PC/MA/EA/DOL/DEM/DMM/LiPF6/LiFSI 3%/21%/22%/10%/15%/19%/4%/6% 17.32 734.3
3 MA/EA/DOL/MTFA/LiPF6/LiFSI 22%/17%/42%/9%/1%/9% 16.75 713.1
4 MA/DEC/EA/EMC/LiFSI 19%/5%/43%/7%/10% 15.95 723.2
5 THF/METHF/EMC/EFA/DME/LiPF6/LiFSI 43%/2%/3%/5%/38%/1%/8% 15.92 711.1
6 MA/EA/EFA/MTFA/LiFSI 48%/3%/5%/34%/10% 15.78 727.2
7 MA/EA/EGDEE/DEM/LiFSI 34%/46%/3%/7%/10% 15.59 717.1
8 EA/DOL/DEM/EP/LiPF6/LiFSI 7%/32%/25%/26%/2%/8% 14.66 712.1
9 EC/EA/THF/EGDEE/METHF/FEC/LiFSI 3%/14%/4%/21%/7%/42%/9% 12.79 724.2
10 MA/EA/DOL/DEM/LiFSI 11%/36%/27%/16%/10% 11.99 719.2
[…]
Figure S1: Detailed model architectures within the workflow reported in this work. a Overall workflow of both predictive and generative process. Solvents and Li salts are considered separately in the workflow. b Module architecture of GNN model which takes SMILES of a single molecule as input and generates a universal molecular embedding by multi-task learning on 11 molecular properties. c. A self-attention-based aggregation block to merge multiple molecular embeddings into a mixture embedding and ensure permutation invariance. $\times$ is row-wise multiplication, $\odot$ stands for matrix multiplication and $\langle,\rangle$ represents inner product. d The empirical relation block for conductivity and anion ratio prediction. For conductivity, there are six empirical learnable parameters (except viscosity) based on the electrolyte embedding, while anion ratio is predicted directly from the electrolyte embedding concatenated with temperature and concentration using a readout MLP layer. Activation layers are omitted from the schematic. e The decoder module which recovers molecular embeddings from mixture embeddings for both solvents and salts, which is further matched to molecules in our electrolyte molecule database.